OpenStudy (18jonea):

Divide (3x3 + 4x2 – 5x – 2) ÷ (x + 2) using long division. 3x2 + 2x + 1 3x2 + 2x – 1 3x2 – 2x – 1 3x2 – 2x + 1 @amistre64

OpenStudy (18jonea):

@Michele_Laino

OpenStudy (18jonea):

@freckles

OpenStudy (18jonea):

@jagr2713

OpenStudy (18jonea):

@amistre64

OpenStudy (18jonea):

@Agl202

OpenStudy (amistre64):

what is your attempt?

OpenStudy (18jonea):

well i know the first term is 3x^2 and then + 2x

OpenStudy (18jonea):

i thinl

OpenStudy (amistre64):

(x+2) (ax^2+bx+c) what is the exapnasion?

OpenStudy (michele_laino):

first step: we have to divide \(3x^3\) by \(x\)

OpenStudy (18jonea):

which is 3x^2

OpenStudy (amistre64):

ax^3 +bx^2 +cx +2ax^2 +2bx +2c ax^3 +(b+2a)x^2 +(c+2b)x +2c 3 4 -5 -2 a=3, c=-1, what is b?

OpenStudy (michele_laino):

yes! so we can write this: 

OpenStudy (18jonea):

-2 is b

OpenStudy (amistre64):

then we have what we need, 3, -2, -1

OpenStudy (amistre64):

it wasnt dividing it of course .. but its a fair way to double chk your results

OpenStudy (18jonea):

ok so the answer is d

OpenStudy (amistre64):

D is 3,-2,1 so no

OpenStudy (18jonea):

sorry i mean c

OpenStudy (amistre64):

unless you have some lettering scheme that differs from top to bottom

OpenStudy (amistre64):

the third one, is 3,-2,-1 should fit the bill yes

OpenStudy (18jonea):

ok thank you i Have one more problem that I have to get don, Could you help me?

OpenStudy (amistre64):

maybe

OpenStudy (18jonea):

Use synthetic division and the Remainder Theorem to find P(–3) if P(x) = x4 + 19x3 + 108x2 + 236x + 176. 9 8 7 6

OpenStudy (18jonea):

after x are exponents

OpenStudy (amistre64):

seems like a sadistic exercise. its simpler to calculate p(-3) directly

OpenStudy (amistre64):

show your efforts please

OpenStudy (18jonea):

I dont even know where to start with this

OpenStudy (amistre64):

if your interest in learning synthetic division ... thats one thing, otherwise just calculate P(-3) directly

OpenStudy (18jonea):

Right so how would I calculate p-(3)

OpenStudy (amistre64):

let x=-3

OpenStudy (amistre64):

synthetic just turns this into a repetition of adding and multiplication ... which essentially makes you work the problem directly in an indirect manner

OpenStudy (amistre64):

spose we had: ax^2+bx+c a b c 0 ka k^2a+kb -------------------- k| a ka+b k^2a +kb +c ^^^^^^^^^^ the final result is ak^2 + bk + c

OpenStudy (18jonea):

i got 8 when x=-3

OpenStudy (amistre64):

so did i

OpenStudy (18jonea):

so would that be the answer to the question

OpenStudy (amistre64):

of course

OpenStudy (18jonea):

ok thank you so much!!!!

OpenStudy (amistre64):

x4 + 19x3 + 108x2 + 236x + 176 by factoring we can break this down into the same synthetic process x(x3 + 19x2 + 108x + 236) + 176 x(x(x2 + 19x + 108) + 236) + 176 x(x(x(x + 19) + 108) + 236) + 176 let x=-3 -3(-3(-3(-3 + 19) + 108) + 236) + 176 -3(-3(-3(16) + 108) + 236) + 176 -3(-3(-48 + 108) + 236) + 176 -3(-3(60) + 236) + 176 -3(-180 + 236) + 176 -3(56) + 176 -168 + 176 -8 + 16 = 8